OpenStudy (anonymous):

the second angle of a triangular garden is four times as large as the first. The third angle is 45 degrees less than the sum of the other two angles. Find the measure of each angle.

OpenStudy (anonymous):

I'm not sure what you said I guest x+4x+45=180 The sum of triangle is 180 degree 5x =180-45 x = 135/5 x =27 4*27=108 the angles is: 45,45,108 I hope it help

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

you are welcome